OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (2.4, 3) and (5, y) 1.44 1 0.69 6.25

OpenStudy (zpupster):

you need to find the constant of variation: y=k/x the first points 3=k/2.4 7.2=k next y=7.2/5 = 1.44so (2.4,1.44)